23g ethanol is dissolved in 36g water.find mole fraction of ethanol?
 

Grade:12th pass

1 Answers

Aniket Raj
2 years ago
no.of moles of ethanol =mass÷molar mass=23÷46=0.5 moles
no.of moles of water=36÷18=2moles
mole fraction of ethanol=0.5÷(0.5+2)=0.2 moles